# Functional Central Limit Theorem for the simultaneous subgraph count of dynamic Erdős-Rényi random graphs

### v1

@@ Appendix A Auxiliary proofs
$\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$ $=\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{1}\}\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{2}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
$=\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{1}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{2}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}.$ (A.8)
We would like to show that for any $g^{\star}\sim g\in{\tt OCS}(H,H^{\star})$
$\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{1}\cap K_{\boldsymbol{v}}\}=\mathbb{I}\{g^{\star}=g_{1}\cap K_{\boldsymbol{v}}\},\qquad\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{2}\cap K_{\boldsymbol{v}}\}=\mathbb{I}\{g^{\star}=g_{2}\cap K_{\boldsymbol{v}}\}.$ (A.9)
In other words, we show that if (A.9) is not satisfied, $g^{\star}$ is not an optimal common subgraph of $H$ and $H^{\star}$. In case $\varrho_{N}=1$, we see from combining (3.8) and (3.3) that the only optimal common subgraph possible is an edge, meaning that $g^{\star}$ is isomorphic to an edge. Hence, as $\left\lvert\boldsymbol{v}\right\rvert=\mathcal{V}(g^{\star})=2$, graph $K_{\boldsymbol{v}}$ as well is isomorphic to an edge, thus (A.9) follows. So, we need to consider only the case $\varrho_{N}\to 0$ as $N\to\infty$. As ${H}\cong{H^{\star}}$ and $g\in{\tt OCS}({H},{H^{\star}})$, applying (3.12) we have that
@@
$g^{\star}{\,\sqsubset\,}g_{1}\cap K_{\boldsymbol{v}},\qquad g^{\star}\not=g_{1}\cap K_{\boldsymbol{v}}.$ (A.10)
Define $g^{\star\star}=g_{1}\cap K_{\boldsymbol{v}}$. It is clear that $g^{\star\star}{\,\sqsubset\,}{\tt CS}({H})$, $\mathcal{V}(g^{\star\star})=\mathcal{V}(g^{\star})=\mathcal{V}(g)$, but due to (A.10) $\mathcal{E}(g^{\star\star})>{\mathcal{E}(g^{\star})=\mathcal{E}(g)}$. Hence,
@@
which contradicts (3.9) as $g^{\star\star}\in{\tt CS}({H})$ and $g\in{\tt OCS}({H})$. This implies that (A.9) is met. Applying it to (A.8) we obtain that
$\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$ $=\mathbb{I}\{g^{\star}=g_{1}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g^{\star}{\,\sqsubset\,}g_{2}=K_{\boldsymbol{v}}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
@@
$\mathcal{C}({H},{H^{\star}},g)\>=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,\sqsubset\,}K_{\boldsymbol{v}}\\ g^{\star}\sim g\end{subarray}}$ $\left(\sum_{\begin{subarray}{c}g_{1}{\,\sqsubset\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}=g^{\star}\}\right)$
[…]
@@
$\sum_{\begin{subarray}{c}g_{1}{\,\sqsubset\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}=g^{\star}\}=\frac{\mathcal{K}({H},g)}{\left\lvert{\tt G}_{\mathcal{V}(g)}(g)\right\rvert}=\frac{\mathcal{K}({H},g)\mathcal{A}(g)}{\mathcal{V}(g)!},$
@@
$\mathcal{K}({H},g)=\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathbb{I}\{g_{1}\cap K_{\langle\mathcal{V}(g)\rangle}\sim g\}.$
@@
$\sum_{\begin{subarray}{c}g_{2}{\,\sqsubset\,}K_{\langle\mathcal{N}\rangle\setminus\boldsymbol{v}_{1}}\\ g_{2}\sim{H^{\star}}\end{subarray}}\mathbb{I}\{g_{2}\cap K_{\boldsymbol{v}}=g^{\star}\}=\frac{\mathcal{K}({H^{\star}},g)\mathcal{A}(g)}{\mathcal{V}(g)!}.$
@@
$\mathcal{C}({H},{H^{\star}},g)$ $=\frac{\mathcal{K}({H},g)\mathcal{K}({H^{\star}},g)\mathcal{A}^{2}(g)}{\bigl{(}\mathcal{V}(g)!\bigr{)}^{2}\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,\sqsubset\,}K_{\boldsymbol{v}}\\ g^{\star}\sim g\end{subarray}}1$
@@
Finally, consider the constants $\mathcal{K}({H},g)$. Let us for all $g_{i}\in{\tt G}_{\mathcal{V}({H})}({H})$ calculate the total number of subgraphs which are isomorphic to $g$. On one hand, this number can be represented using the constant $\mathcal{K}({H},g)$:
$\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))}\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}\sim g\}=\left\lvert{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))\right\rvert\mathcal{K}({H},g).$
@@
$\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}\sim g\}$ $=\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathcal{S}({H},g)=\left\lvert{\tt G}_{\mathcal{V}({H})}({H})\right\rvert\mathcal{S}({H},g).$
@@
Proof of Proposition 5.3. Using the definition of process $\boldsymbol{X}(\cdot)$ given in Theorem 4.1,
@@
×$\qquad\qquad\qquad\qquad\qquad\times\mathcal{S}({H}_{i},g_{1})\,\mathcal{S}({H}_{j},g_{2})\operatorname{Cov}(X_{g_{1}}^{\star}(t),X^{\star}_{g_{2}}(s))$
@@
×$\qquad\qquad\qquad\qquad\qquad\times\mathcal{S}({H}_{i},{g})\,\mathcal{S}({H}_{j},{g})\bigl{(}\kappa_{\star}(t,s)\bigr{)}^{\mathcal{E}(g)}.$
where in the last line we use the independence of the individual processes $X^{\star}_{g}(\cdot)$. Hence, the claim follows from applying (3.12), (3.13) and Lemma 3.4. $\Box$
Proof of Lemma 8.1. For every edge $(u,v)\in{\tt E}(K_{N})$ and time $t^{\star}\in[0,T]$, define the event that the edge $(u,v)$ exists at time $t^{\star}$:
@@
$\Omega_{k}=\left(\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}(x_{k})\right)\cup\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}(y_{k})\right)\right){\bigcap}\left(\bigcup_{(u,v)\in{\tt E}(g_{k})}\Omega^{\not=}_{(u,v)}(x_{k},y_{k})\right)$ (A.15)
@@
$\Omega_{k}$ $\subset\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}\right){\bigcap}\left(\bigcup_{(u,v)\in{\tt E}(g_{k})}\Omega^{\not=}_{(u,v)}(x_{k},y_{k})\right)$ (A.16)
@@
$\mathbb{P}\left\{\Omega^{+}_{(u,v)}\right\}$ $\leqslant{p_{+}}\,\varrho_{N}.$ (A.20)
@@
$\mathbb{P}\left\{\Omega_{k}\right\}$ $\leqslant\sum_{{(u^{\star},v^{\star})\in{\tt E}(g_{k})}}\mathbb{P}\left\{\Omega^{\not=}_{(u^{\star},v^{\star})}(x_{k},y_{k})\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}\right)\right\}$
@@
$\mathbb{P}\left\{\Omega_{k_{1}}\Omega_{k_{2}}\Omega_{k_{3}}\right\}$ $\leqslant\sum_{\begin{subarray}{c}(u_{1}^{\star},v_{1}^{\star})\in{\tt E}(g_{k_{1}})\\ (u_{2}^{\star},v_{2}^{\star})\in{\tt E}(g_{k_{2}})\end{subarray}}\mathbb{P}\left\{\Omega^{\not=}_{(u_{1}^{\star},v_{1}^{\star})}(r,s)\Omega^{\not=}_{(u_{2}^{\star},v_{2}^{\star})}(s,t)\left(\bigcap_{(u,v)\in{\tt E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{3})}\Omega^{+}_{u,v}\right)\right\}$
@@
$\leqslant C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}})}\left\lvert t-r\right\rvert^{2},$
@@
$\mathbb{P}\left\{\Omega_{k_{1}}\Omega_{k_{2}}\Omega_{k_{3}}\Omega_{k_{4}}\right\}$ $\leqslant{\mathfrak{C}^{2}}\,\mathcal{E}(g_{k_{1}})\mathcal{E}(g_{k_{2}})\,{p_{+}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{4})-2}}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{4})}\left\lvert t-r\right\rvert^{2}$
$\qquad+\,{\mathfrak{C}}\,\mathcal{E}(g_{k_{1}}\cap g_{k_{2}})\,{p_{+}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{4})-1}}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{4})}\left\lvert t-r\right\rvert^{2}$
$\leqslant C_{0}\,\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{4})}\left\lvert t-r\right\rvert^{2},$
Hence, the claim of Lemma 8.1 follows. $\Box$
Proof of Lemma 7.1. We establish the statement of this lemma for any connected graph ${\tt G}{\,\sqsubset\,}K_{v}$ with at least two vertices; recall that from (7.9) we know that $\mathcal{V}({\tt G})\geqslant 2$ for any ${\tt G}\in{\tt CC}(\mathcal{G}(\boldsymbol{g}))$. As $N\to\infty$,
$\mathbb{E}\left\{N^{\mathcal{V}({\tt G})}\prod_{i\in{\tt V}({\tt G})}\frac{\mathcal{F}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\right\}$ ${=N^{\mathcal{V({\tt g})}}\mathbb{E}\left\{\prod_{i\in{\tt V}({\tt g})}\xi_{f_{i},N}(g_{i})\right\}\prod_{i\in{\tt V}({\tt G})}\frac{\mathcal{F}_{N}({H}_{f_{i}})}{\sqrt{\mathcal{F}_{N}^{\rm{opt}}({H}_{f_{i}})}}=:\mathcal{E}_{N}({\tt G})}$ (A.23)
@@
$\prod_{i\in{\tt V}({\tt g})}\xi_{f_{i},N}(g_{i})=\sum_{k_{1},\ldots,k_{\mathcal{V}({\tt g})}=1}^{n}\prod_{i\in{\tt V}({\tt g})}\frac{q_{k,f_{i}}\bigl{(}\mathbb{I}\{g_{i}{\,\sqsubset\,}G_{N}(t_{k_{i}})\}-\mathbb{P}\left\{g_{i}{\,\sqsubset\,}G_{N}(t_{k_{i}})\right\}\bigr{)}}{\bigl{(}p_{\star}(t_{k_{i}})\bigr{)}^{\mathcal{E}(g_{i})}}$ (A.24)
@@
$\mathbb{E}\left\{\prod_{i\in{\tt V}({\tt g})}\bigl{(}\mathbb{I}\{g_{i}{\,\sqsubset\,}G_{N}(t_{k_{i}})\}-\mathbb{P}\left\{g_{i}{\,\sqsubset\,}G_{N}(t_{k_{i}})\right\}\bigr{)}\right\}\leqslant(p_{+}\varrho_{N})^{\mathcal{E}(\cup_{i\in{\tt V}({\tt G})}g_{i})}.$ (A.25)
@@
$=:\bigl{(}C_{{\tt G}}+o(1)\bigr{)}\mathcal{E}_{N}({\tt G})$
So, to verify the statement of Lemma 7.1 it is enough to show that
@@
$=\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\sqrt{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}\sqrt{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}=\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}$
By virtue of (3.10) (as $(g_{i}\cap g_{j}){\,\sqsubset\,}g_{i}\sim{H}_{f_{i}}$ and $(g_{i}\cap g_{j}){\,\sqsubset\,}g_{j}\sim{H}_{f_{j}}$), we obtain that
@@
and (A.28) follows. We now proceed to the case $\mathcal{V}({\tt G})\geqslant 3$. We are going to show (A.29) by induction to $\mathcal{V}({\tt G})$. Our induction step is the following. Assume for any graph ${\tt G}{\,\sqsubset\,}K_{v}$ such that $\mathcal{V}({\tt G})=k-1$ that the statement (A.28) applies. Then for any ${\tt G}{\,\sqsubset\,}K_{v}$ such that $\mathcal{V}({\tt G})=k$ we find (A.29).
For any connected graph ${\tt G}{\,\sqsubset\,}K_{v}$ with $k$ vertices, we can find a vertex $j\in{\tt V}({\tt G})$ such that ${\tt G}^{\star}={\tt G}\setminus\{j\}$ is still a connected graph. Then,

### v2

@@ Appendix A Auxiliary proofs
$\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$ $=\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{1}\}\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{2}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
$=\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{1}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{2}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}.$ (A.8)
We would like to show that for any $g^{\star}\subset K_{\boldsymbol{v}}$ such that $\pi(g^{\star})\in{\tt OCS}(H,H^{\star})$
$\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{1}\cap K_{\boldsymbol{v}}\}=\mathbb{I}\{g^{\star}=g_{1}\cap K_{\boldsymbol{v}}\},\qquad\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{2}\cap K_{\boldsymbol{v}}\}=\mathbb{I}\{g^{\star}=g_{2}\cap K_{\boldsymbol{v}}\}.$ (A.9)
In other words, we show that if (A.9) is not satisfied, the pattern of the graph $g^{\star}$ is not an optimal common subgraph pattern of $H$ and $H^{\star}$. In case $\varrho_{N}=1$, we see from combining (3.7) and (3.4) that the only optimal common subgraph possible is an edge, meaning that $g^{\star}$ is isomorphic to an edge. Hence, as $\left\lvert\boldsymbol{v}\right\rvert=\mathcal{V}(g^{\star})=2$, graph $K_{\boldsymbol{v}}$ as well is isomorphic to an edge, thus (A.9) follows. So, we need to consider only the case $\varrho_{N}\to 0$ as $N\to\infty$. As ${H}\cong{H^{\star}}$ and $g\in{\tt OCS}({H},{H^{\star}})$, applying (3.11) we have that
@@
$g^{\star}{\,{\subset}\,}g_{1}\cap K_{\boldsymbol{v}},\qquad g^{\star}\not=g_{1}\cap K_{\boldsymbol{v}}.$ (A.10)
Define $g^{\star\star}=g_{1}\cap K_{\boldsymbol{v}}$. It is clear that ${\pi(g^{\star\star})\in\,}{\tt CS}({H})$, $\mathcal{V}(g^{\star\star})=\mathcal{V}(g^{\star})=\mathcal{V}(g)$, but due to (A.10) $\mathcal{E}(g^{\star\star})>{\mathcal{E}(g^{\star})=\mathcal{E}(g)}$. Hence,
@@
which contradicts (3.8) as ${\pi(g^{\star\star})}\in{\tt CS}({H})$ and $g\in{\tt OCS}({H})$. This implies that (A.9) is met. Applying it to (A.8) we obtain that
$\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$ $=\mathbb{I}\{g^{\star}=g_{1}\cap K_{\boldsymbol{v}}\}\mathbb{I}\{g^{\star}{\,{\subset}\,}g_{2}=K_{\boldsymbol{v}}\}\mathbb{I}\{g_{1}\cap g_{2}=g^{\star}\}$
@@
$\mathcal{C}({H},{H^{\star}},g)\>=\frac{1}{\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,{\subset}\,}K_{\boldsymbol{v}}\\ {\pi(g^{\star})=g}\end{subarray}}$ $\left(\sum_{\begin{subarray}{c}g_{1}{\,{\subset}\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}=g^{\star}\}\right)$
$\hskip 179.25244pt\times\,$ $\left(\sum_{\begin{subarray}{c}g_{2}{\,{\subset}\,}K_{\langle\mathcal{N}\rangle\setminus\boldsymbol{v}_{1}}\\ g_{2}\sim{H^{\star}}\end{subarray}}\mathbb{I}\{g_{2}\cap K_{\boldsymbol{v}}=g^{\star}\}\right){.}$ (A.12)
@@
$\sum_{\begin{subarray}{c}g_{1}{\,{\subset}\,}K_{\boldsymbol{v}\cup\boldsymbol{v}_{1}}\\ g_{1}\sim{H}\end{subarray}}\mathbb{I}\{g_{1}\cap K_{\boldsymbol{v}}=g^{\star}\}=\frac{\mathcal{K}({H},g)}{\left\lvert{\tt G}_{\mathcal{V}(g)}(g)\right\rvert}=\frac{\mathcal{K}({H},g)\mathcal{A}(g)}{\mathcal{V}(g)!},$
@@
$\mathcal{K}({H},g)=\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathbb{I}\{{\pi(g_{1}\cap K_{\langle\mathcal{V}(g)\rangle})=g}\}.$
@@
$\sum_{\begin{subarray}{c}g_{2}{\,{\subset}\,}K_{\langle\mathcal{N}\rangle\setminus\boldsymbol{v}_{1}}\\ g_{2}\sim{H^{\star}}\end{subarray}}\mathbb{I}\{g_{2}\cap K_{\boldsymbol{v}}=g^{\star}\}=\frac{\mathcal{K}({H^{\star}},g)\mathcal{A}(g)}{\mathcal{V}(g)!}.$
@@
$\mathcal{C}({H},{H^{\star}},g)$ $=\frac{\mathcal{K}({H},g)\mathcal{K}({H^{\star}},g)\mathcal{A}^{2}(g)}{\bigl(\mathcal{V}(g)!\bigr)^{2}\mathcal{N}!}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{N}\rangle}(\mathcal{V}(g))}\sum_{\boldsymbol{v}_{1}\in{\tt P}_{\langle N\rangle\setminus\boldsymbol{v}}(\mathcal{V}({H})-\mathcal{V}(g))}\sum_{\begin{subarray}{c}g^{\star}{\,{\subset}\,}K_{\boldsymbol{v}}\\ {\pi(g^{\star})=g}\end{subarray}}1$
@@
Finally, consider the constants $\mathcal{K}({H},g)$. Let us for all $g_{i}\in{\tt G}_{\mathcal{V}({H})}({H})$ calculate the total number of subgraphs which have pattern $g$. On one hand, this number can be represented using the constant $\mathcal{K}({H},g)$:
$\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))}\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathbb{I}\{{\pi(g_{1}\cap K_{\boldsymbol{v}})=g}\}=\left\lvert{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))\right\rvert\mathcal{K}({H},g).$
@@
$\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\sum_{\boldsymbol{v}\in{\tt P}_{\langle\mathcal{V}({H})\rangle}(\mathcal{V}(g))}\mathbb{I}\{{\pi(g_{1}\cap K_{\boldsymbol{v}})=g}\}$ $=\sum_{g_{1}\in{\tt G}_{\mathcal{V}({H})}({H})}\mathcal{S}({H},g)=\left\lvert{\tt G}_{\mathcal{V}({H})}({H})\right\rvert\mathcal{S}({H},g).$
@@
Proof of Proposition 4. Using the definition of process $\boldsymbol{X}(\cdot)$ given in Theorem 4.1,
@@
$\qquad\qquad\qquad\qquad\qquad\times\mathcal{S}({H}_{i},g_{1})\,\mathcal{S}({H}_{j},g_{2})\operatorname{Cov}(X_{g_{1}}^{\star}(t),X^{\star}_{g_{2}}(s))$
@@
$\qquad\qquad\qquad\qquad\qquad\times\mathcal{S}({H}_{i},{g})\,\mathcal{S}({H}_{j},{g})\bigl(\kappa_{\star}(t,s)\bigr)^{\mathcal{E}(g)}.$
where in the last line we use the independence of the individual processes $X^{\star}_{g}(\cdot)$. Hence, the claim follows from applying (3.11), (3.12) and Lemma 2. $\Box$
Proof of Lemma 9. For every edge $(u,v)\in{\tt E}(K_{N})$ and time $t^{\star}\in[0,T]$, define the event that the edge $(u,v)$ exists at time $t^{\star}$:
@@
$\Omega_{k}=\left(\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}(x_{k})\right)\cup\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}(y_{k})\right)\right){\bigcap}\left(\bigcup_{(u,v)\in{\tt E}(g_{k})}\Omega^{\not=}_{(u,v)}(x_{k},y_{k})\right){.}$ (A.15)
@@
$\Omega_{k}$ $\subset\left(\bigcap_{(u,v)\in{\tt E}(g_{k})}\Omega^{+}_{{(u,v)}}\right){\bigcap}\left(\bigcup_{(u,v)\in{\tt E}(g_{k})}\Omega^{\not=}_{(u,v)}(x_{k},y_{k})\right){.}$ (A.16)
@@
$\mathbb{P}\left\{\Omega^{+}_{(u,v)}\right\}$ $\leqslant{p_{+}}\,\varrho_{N}{,}$ (A.20)
@@
[…]
@@
Applying now all three bounds (A.20), (A.21) and (A.22) to the probability of the event (A.18), we obtain
@@
$\leqslant C_{0}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}})}\left\lvert t-r\right\rvert^{2}{.}$
@@
$\mathbb{P}\left\{\Omega_{k_{1}}\Omega_{k_{2}}\Omega_{k_{3}}\Omega_{k_{4}}\right\}$ $\leqslant{\mathfrak{C}^{2}}\,\mathcal{E}(g_{k_{1}})\mathcal{E}(g_{k_{2}})\,{p_{+}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{k_{4}})-2}}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{k_{4}})}\left\lvert t-r\right\rvert^{2}$
$\qquad+\,{\mathfrak{C}}\,\mathcal{E}(g_{k_{1}}\cap g_{k_{2}})\,{p_{+}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{k_{4}})-1}}\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{k_{4}})}\left\lvert t-r\right\rvert^{2}$
$\leqslant C_{0}\,\varrho_{N}^{\mathcal{E}(g_{k_{1}}\cup g_{k_{2}}\cup g_{k_{3}}\cup g_{k_{4}})}\left\lvert t-r\right\rvert^{2}{.}$
Hence, the claim of Lemma 9 follows. $\Box$
Proof of Lemma 7. We establish the statement of this lemma for any connected graph ${\tt G}{\,{\subset}\,}K_{v}$ with at least two vertices; recall that from (7.9) we know that $\mathcal{V}({\tt G})\geqslant 2$ for any ${\tt G}\in{\tt CC}(\mathcal{G}(\boldsymbol{g}))$. As $N\to\infty$,
$\mathbb{E}\left\{N^{\mathcal{V}({\tt G})}\prod_{i\in{\tt V}({\tt G})}\frac{\mathcal{F}({H}_{f_{i}})\xi_{f_{i},{N}}(g_{i})}{\sqrt{\mathcal{F}^{\rm opt}_{N}({H}_{f_{i}})}}\right\}$ ${=N^{\mathcal{V({\tt g})}}\mathbb{E}\left\{\prod_{i\in{\tt V}({\tt g})}\xi_{f_{i},N}(g_{i})\right\}\prod_{i\in{\tt V}({\tt G})}\frac{\mathcal{F}_{N}({H}_{f_{i}})}{\sqrt{\mathcal{F}_{N}^{\rm{opt}}({H}_{f_{i}})}}=:\mathcal{E}_{N}({\tt G})}{.}$ (A.23)
@@
$\prod_{i\in{\tt V}({\tt g})}\xi_{f_{i},N}(g_{i})=\sum_{k_{1},\ldots,k_{\mathcal{V}({\tt g})}=1}^{n}\prod_{i\in{\tt V}({\tt g})}\frac{q_{k,f_{i}}\bigl(\mathbb{I}\{g_{i}{\,{\subset}\,}G_{N}(t_{k_{i}})\}-\mathbb{P}\left\{g_{i}{\,{\subset}\,}G_{N}(t_{k_{i}})\right\}\bigr)}{\bigl(p_{\star}(t_{k_{i}})\bigr)^{\mathcal{E}(g_{i})}}{.}$ (A.24)
@@
$\mathbb{E}\left\{\prod_{i\in{\tt V}({\tt g})}\bigl(\mathbb{I}\{g_{i}{\,{\subset}\,}G_{N}(t_{k_{i}})\}-\mathbb{P}\left\{g_{i}{\,{\subset}\,}G_{N}(t_{k_{i}})\right\}\bigr)\right\}\leqslant(p_{+}\varrho_{N})^{\mathcal{E}(\cup_{i\in{\tt V}({\tt G})}g_{i})}.$ (A.25)
@@
$=:\bigl(C_{{\tt G}}+o(1)\bigr)\mathcal{E}_{N}({\tt G}){.}$
So, to verify the statement of Lemma 7 it is enough to show that
@@
$=\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\sqrt{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}\sqrt{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}=\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{i}})}}\sqrt{\frac{\mathcal{F}_{N}(g_{i}\cap g_{j})}{\mathcal{F}^{\rm{opt}}_{N}({H}_{f_{j}})}}{.}$
By virtue of (3.9) (as $(g_{i}\cap g_{j}){\,{\subset}\,}g_{i}\sim{H}_{f_{i}}$ and $(g_{i}\cap g_{j}){\,{\subset}\,}g_{j}\sim{H}_{f_{j}}$), we obtain that
@@
and (A.28) follows. We now proceed to the case $\mathcal{V}({\tt G})\geqslant 3$. We are going to show (A.29) by induction to $\mathcal{V}({\tt G})$. Our induction step is the following. Assume for any graph ${\tt G}{\,{\subset}\,}K_{v}$ such that $\mathcal{V}({\tt G})=k-1$ that the statement (A.28) applies. Then for any ${\tt G}{\,{\subset}\,}K_{v}$ such that $\mathcal{V}({\tt G})=k$ we find (A.29).
For any connected graph ${\tt G}{\,{\subset}\,}K_{v}$ with $k$ vertices, we can find a vertex $j\in{\tt V}({\tt G})$ such that ${\tt G}^{\star}={\tt G}\setminus\{j\}$ is still a connected graph. Then,